# VisJudge-Bench: Aesthetics and Quality Assessment of Visualizations

### v1

@@ Abstract
Visualization, a domain-specific yet widely used form of imagery, is an effective way to turn complex datasets into intuitive insights, and its value depends on whether data are faithfully represented, clearly communicated, and aesthetically designed. However, evaluating visualization quality is challenging: unlike natural images, it requires simultaneous judgment across data encoding accuracy, information expressiveness, and visual aesthetics. Although multimodal large language models (MLLMs) have shown promising performance in aesthetic assessment of natural images, no systematic benchmark exists for measuring their capabilities in evaluating visualizations. To address this, we propose VisJudge-Bench, the first comprehensive benchmark for evaluating MLLMs’ performance in assessing visualization aesthetics and quality. It contains 3,090 expert-annotated samples from real-world scenarios, covering single visualizations, multiple visualizations, and dashboards across 32 chart types. Systematic testing on this benchmark reveals that even the most advanced MLLMs (such as GPT-5) still exhibit significant gaps compared to human experts in judgment, with a Mean Absolute Error (MAE) of 0.551 and a correlation with human ratings of only 0.429. To address this issue, we propose VisJudge, a model specifically designed for visualization aesthetics and quality assessment. Experimental results demonstrate that VisJudge significantly narrows the gap with human judgment, reducing the MAE to 0.442 (a 19.8% reduction) and increasing the consistency with human experts to 0.681 (a 58.7% improvement) compared to GPT-5. The benchmark is available at https://github.com/HKUSTDial/VisJudgeBench.
@@ 1 Introduction
Visualization serves as an effective approach for transforming complex datasets into intuitive insights (Shen et al., 2023; Qin et al., 2020; Ye et al., 2024; Qin et al., 2020; Li et al., 2024a; 2025). The value of a high-quality visualization depends on whether its data is faithfully presented (Fidelity), whether information is clearly communicated (Expressiveness), and whether the design is aesthetically well-presented (Aesthetics), as shown in Figure 1. These three dimensions are closely interconnected and indispensable, posing challenges for visualization quality assessment.
@@
To address this challenge, we construct VisJudge-Bench, the first comprehensive benchmark based on the “Fidelity, Expressiveness, and Aesthetics” principles to assess MLLMs’ capabilities in visualization aesthetics and quality evaluation. It contains 3,090 expert-scored samples from real-world scenarios, covering single visualizations, multiple visualizations, and dashboards across 32 chart types. Using this benchmark, we conduct extensive testing on multiple MLLMs, including GPT-5, finding that even the most advanced models show significant differences from human experts (Mean Absolute Error as high as 0.551, correlation only 0.429). This finding clearly demonstrates that general MLLMs cannot automatically acquire specialized evaluation capabilities in the visualization domain, making the development of specialized optimization models necessary.
Based on this, we propose VisJudge, a model specifically designed for visualization aesthetics and quality assessment, aimed at improving the consistency between general MLLMs and human expert evaluation standards. Experimental results prove the effectiveness of this approach: VisJudge significantly improves consistency with human experts, achieving a 19.8% reduction in MAE (to 0.442) and a 58.7% improvement in correlation (to 0.681) compared to GPT-5, performing best among all tested models.
@@ 3.1.1 Data Collection and Preprocessing
Corpus Construction. To evaluate the performance of MLLMs across different visualization types, we construct a corpus covering three main categories: single visualizations, multiple visualizations, and dashboards. To ensure the authenticity and diversity of our corpus, we collect visualization samples from search engines using web crawling methods with diverse query keywords (see Appendix A.1.1 for detailed crawling architecture and keyword generation strategy).
Data Preprocessing Pipeline. We design a three-stage data filtering process to curate the benchmark from over 300,000 initial images. (1) Initial Filtering: We employ automated scripts and perceptual hash algorithms to eliminate non-visualization content and duplicates, yielding 80,210 candidate images (detailed algorithms in Appendix A.1.1). (2) Automated Classification: We leverage GPT-4o for visualization type classification and quality filtering, resulting in 13,220 valid visualization samples after human verification (classification prompts and criteria in Appendix A.1.1). (3) Stratified Sampling: We apply stratified random sampling to select the final 3,090 samples, ensuring balanced distribution across categories. As shown in Table 2, the final corpus contains 1,041 single visualizations, 1,024 multiple visualizations, and 1,025 dashboards, covering 32 distinct subtypes. Complete statistical breakdown is provided in Appendix A.2.1.
@@ 3.1.2 The “Fidelity, Expressiveness, and Aesthetics” Evaluation Framework
The “Fidelity, Expressiveness, and Aesthetics” Framework Design. To systematically evaluate visualization quality, we construct a multi-dimensional evaluation framework. This framework draws inspiration from classical translation theory principles of “Fidelity, Expressiveness, and Aesthetics” (illustrated with positive-negative examples in Figure 1), combined with established theories in graphical perception (Cleveland & McGill, 1984), information visualization design (Munzner, 2014), and aesthetic evaluation (Li et al., 2024c). We operationalize this core concept into six measurable evaluation dimensions (as shown in Figure 3):
  • Fidelity focuses on Data Fidelity. This dimension draws from Tufte’s design principles (Tufte, 1983) for avoiding “graphical lies” and recent research on visualization misleadingness issues (Nguyen et al., 2013; Szafir, 2018; Lan & Liu, 2024; McNutt et al., 2020). It primarily evaluates whether visual encodings accurately reflect the original data, avoiding misleading interpretations caused by improper axis settings, scale distortions, or other design flaws.
  • Expressiveness focuses on the effectiveness of information communication. This dimension evaluates how effectively visualizations convey information to users. It includes two progressive sub-dimensions: First, (1) Semantic Readability evaluates the clarity of basic information encoding, assessing whether users can unambiguously decode visual elements in charts (Pan et al., 2025). Building on chart readability, (2) Insight Discovery further evaluates the analytical value in revealing deep data patterns, trends, or outliers, helping users transition from “reading information” to “gaining insights” (Wu et al., 2024a).
  • Aesthetics focuses on Aesthetic Quality of visual design, integrating visualization perception theory (Ware, 2021) with design practice. This dimension consists of three sub-dimensions that collectively influence the overall visual experience: (1) Design Style evaluates the innovation and uniqueness of design, measuring the degree of novel visual elements and distinctive style (Dibia, 2023; Brath & Banissi, 2016); (2) Visual Composition focuses on the rationality of spatial layout, evaluating the balance and order of element positioning, size proportions, and spacing arrangements (Wu et al., 2023); and (3) Color Harmony evaluates the coordination and functionality of color combinations, ensuring color palette choices balance aesthetics with effective information communication (Harrower & Brewer, 2003; Gramazio et al., 2017).
@@
Adaptive Question Generation Mechanism. Based on the evaluation framework, we have devised an adaptive question generation process (detailed workflow shown in Figure 3). This process begins by leveraging GPT-4o to extract metadata from the chart, such as its type and visual elements. Subsequently, it rewrites questions by populating predefined templates based on this metadata, generating highly customized questions for the six evaluation sub-dimensions. This approach ensures that the evaluation questions are closely aligned with the specific visualization content. For more detailed examples, please refer to Appendix C.1.
@@ 3.1.3 Expert Annotation and Quality Control
Stage 1: Initial Annotation. We recruited highly qualified crowdsourcing workers through the CloudResearch platform (CloudResearch, 2022). To ensure annotation quality, we not only set strict screening criteria (e.g., Bachelor’s degree or higher, 97%+ task approval rate, native English speaker, and professional background in multiple relevant fields; see Appendix A.3.1 for details), but also designed a dedicated annotation interface (see Appendix A.3.3) to guide the process. Crucially, we embedded “validation checks” into the annotation tasks to identify and filter out inattentive responses (examples in Appendix A.3.4). We paid USD \$10 per hour for this task. Each of the 3,090 samples was scored by three independent annotators across six evaluation dimensions (see Appendix A.3.2 for task design details), generating an initial scoring matrix.
Stage 2: Quality Control. To address scoring disagreements among annotators, we designed a systematic conflict identification and resolution mechanism based on established crowdsourcing quality control and statistical evaluation theory (Gadiraju et al., 2015; Rousseeuw & Leroy, 2005; Brennan, 2001). The system first identifies high-disagreement samples by analyzing score variance, then algorithmically generates candidate resolution strategies including outlier removal, malicious scoring detection, and sub-dimensional bias correction. These algorithm-generated suggestions are processed and ranked before being submitted to the expert team for final review (complete algorithmic details and parameters in Appendix A.3.4).
Stage 3: Expert Validation. The final review phase is handled by a team of three experts with visualization analysis experience (expert interface and workflow described in Appendix A.3.5). The experts independently review disputed samples and algorithm-generated candidate solutions, using their professional knowledge to select, modify, or reject strategies. For particularly complex disputed samples, the expert team reaches consensus through discussion. Through this end-to-end rigorous process, we ultimately built a high-quality human scoring benchmark for all 3,090 samples, serving as the gold standard for evaluating model performance.
@@ 4 VisJudge: A Specialized Model for Visualization Evaluation
To validate VisJudge-Bench as an effective training resource, we fine-tuned a specialized model called VisJudge to enhance MLLM visualization evaluation capabilities.
@@
Model Training. We selected Qwen2.5-VL-7B-Instruct (Bai et al., 2025) as our base model for its strong multimodal capabilities. The model generates quality scores (1.0-5.0) and rationales aligned with human expert judgments. We employ reinforcement learning with the GRPO algorithm (Shao et al., 2024), using a composite reward function combining accuracy reward (minimizing prediction error) and format reward (ensuring structured outputs) (Shi et al., 2025; Wu et al., 2025). Formal reward definitions are detailed in Appendix D.2. For parameter-efficient fine-tuning, we adopted the Low-Rank Adaptation (LoRA) (Hu et al., 2022). Training used four NVIDIA A6000 (48GB) GPUs for 5 epochs with a learning rate of 1e-5. Detailed configurations are in Appendix D.3.
@@ 5.1 Experimental Settings
Evaluation Setup. We evaluate seven representative MLLMs: GPT-5, GPT-4o, Claude-4-Sonnet, Claude-3.5-Sonnet, Gemini-2.0-Flash, Gemini-2.5-Pro, Qwen2.5-VL-7B-Instruct (Bai et al., 2025), and VisJudge on a balanced test set of 648 samples (see Appendix A.2.3 for distribution details). Each model provides 1-to-5 scores with justifications based on our “Fidelity, Expressiveness, and Aesthetics” framework. Following human annotation procedures, we run each model three times and average the results. All inference uses vLLM on four NVIDIA A6000 (48GB) GPUs with bfloat16 precision and a temperature of 0.8.

### v2

@@ Abstract
Visualization, a domain-specific yet widely used form of imagery, is an effective way to turn complex datasets into intuitive insights, and its value depends on whether data are faithfully represented, clearly communicated, and aesthetically designed. However, evaluating visualization quality is challenging: unlike natural images, it requires simultaneous judgment across data encoding accuracy, information expressiveness, and visual aesthetics. Although multimodal large language models (MLLMs) have shown promising performance in aesthetic assessment of natural images, no systematic benchmark exists for measuring their capabilities in evaluating visualizations. To address this, we propose VisJudge-Bench, the first comprehensive benchmark for evaluating MLLMs’ performance in assessing visualization aesthetics and quality. It contains 3,090 expert-annotated samples from real-world scenarios, covering single visualizations, multiple visualizations, and dashboards across 32 chart types. Systematic testing on this benchmark reveals that even the most advanced MLLMs (such as GPT-5) still exhibit significant gaps compared to human experts in judgment, with a Mean Absolute Error (MAE) of 0.553 and a correlation with human ratings of only 0.428. To address this issue, we propose VisJudge, a model specifically designed for visualization aesthetics and quality assessment. Experimental results demonstrate that VisJudge significantly narrows the gap with human judgment, reducing the MAE to 0.421 (a 23.9% reduction) and increasing the consistency with human experts to 0.687 (a 60.5% improvement) compared to GPT-5. The benchmark is available at https://github.com/HKUSTDial/VisJudgeBench.
@@ 1 Introduction
Visualization serves as an effective approach for transforming complex datasets into intuitive insights (Shen et al., 2023; Qin et al., 2020; Ye et al., 2024; Li et al., 2024a; 2025a). The value of a high-quality visualization depends on whether its data is faithfully presented (Fidelity), whether information is clearly communicated (Expressiveness), and whether the design is aesthetically well-presented (Aesthetics), as shown in Figure 1. These three dimensions are closely interconnected and indispensable, posing challenges for visualization quality assessment.
@@
To address this challenge, we construct VisJudge-Bench, the first comprehensive benchmark based on the “Fidelity, Expressiveness, and Aesthetics” principles to assess MLLMs’ capabilities in visualization aesthetics and quality evaluation. It contains 3,090 expert-scored samples from real-world scenarios, covering single visualizations, multiple visualizations, and dashboards across 32 chart types. Using this benchmark, we conduct extensive testing on multiple MLLMs, including GPT-5, finding that even the most advanced models show significant differences from human experts (Mean Absolute Error as high as 0.553, correlation only 0.428). This finding clearly demonstrates that general MLLMs cannot automatically acquire specialized evaluation capabilities in the visualization domain, making the development of specialized optimization models necessary.
Based on this, we propose VisJudge, a model specifically designed for visualization aesthetics and quality assessment, aimed at improving the consistency between general MLLMs and human expert evaluation standards. Experimental results prove the effectiveness of this approach: VisJudge significantly improves consistency with human experts, achieving a 23.9% reduction in MAE (to 0.421) and a 60.5% improvement in correlation (to 0.687) compared to GPT-5, performing best among all tested models.
@@ 3.1.1 Data Collection and Preprocessing
Corpus Construction. To evaluate the performance of MLLMs across different visualization types, we construct a corpus covering three main categories: single visualizations, multiple visualizations, and dashboards. To ensure the authenticity and diversity of our corpus, we collect visualization samples from search engines using web crawling methods with diverse query keywords (see Appendix A.1 for detailed crawling architecture and keyword generation strategy).
Data Preprocessing Pipeline. We design a three-stage data filtering process to curate the benchmark from over 300,000 initial images. (1) Initial Filtering: We employ automated scripts and perceptual hash algorithms to eliminate non-visualization content and duplicates, yielding 80,210 candidate images (detailed algorithms in Appendix A.1). (2) Automated Classification: We leverage GPT-4o for visualization type classification and quality filtering, resulting in 13,220 valid visualization samples after human verification (classification prompts and criteria in Appendix A.1). (3) Stratified Sampling: We apply stratified random sampling to select the final 3,090 samples, ensuring balanced distribution across categories. As shown in Table 2, the final corpus contains 1,041 single visualizations, 1,024 multiple visualizations, and 1,025 dashboards, covering 32 distinct subtypes. Detailed dataset statistics and distribution analyses are provided in Appendix A.2.
@@ 3.1.2 The “Fidelity, Expressiveness, and Aesthetics” Evaluation Framework
The “Fidelity, Expressiveness, and Aesthetics” Framework Design. To systematically evaluate visualization quality, we construct a multi-dimensional evaluation framework. This framework draws inspiration from classical translation theory principles of “Fidelity, Expressiveness, and Aesthetics” (illustrated with positive-negative examples in Figure 1), combined with established theories in graphical perception (Cleveland and McGill, 1984), information visualization design (Munzner, 2014), and aesthetic evaluation (Li et al., 2024c). We operationalize this core concept into six measurable evaluation dimensions (as shown in Figure 3):
  • Fidelity focuses on Data Fidelity. This dimension draws from Tufte’s design principles (Tufte, 1983) for avoiding “graphical lies” and recent research on visualization misleadingness issues (Nguyen et al., 2013; Szafir, 2018; Lan and Liu, 2024b; McNutt et al., 2020). Since source data for web-collected visualizations is typically unavailable, this dimension evaluates data presentation accuracy at the visual level. It assesses whether visual encodings accurately reflect the displayed values, examining visually detectable distortions such as improper axis settings, scale distortions, truncated baselines, disproportional encodings, or other misleading design patterns.
  • Expressiveness focuses on the effectiveness of information communication. This dimension evaluates how effectively visualizations convey information to users. It includes two progressive sub-dimensions: First, (1) Semantic Readability evaluates the clarity of basic information encoding, assessing whether users can unambiguously decode visual elements in charts (Pan et al., 2025; Tang et al., 2026). Building on chart readability, (2) Insight Discovery further evaluates the analytical value in revealing deep data patterns, trends, or outliers, helping users transition from “reading information” to “gaining insights” (Wu et al., 2024a; Chen et al., 2025).
  • Aesthetics focuses on Aesthetic Quality of visual design, integrating visualization perception theory (Ware, 2021) with design practice. This dimension consists of three sub-dimensions that collectively influence the overall visual experience: (1) Design Style evaluates the innovation and uniqueness of design, measuring the degree of novel visual elements and distinctive style (Dibia, 2023; Brath and Banissi, 2016); (2) Visual Composition focuses on the rationality of spatial layout, evaluating the balance and order of element positioning, size proportions, and spacing arrangements (Wu et al., 2023); and (3) Color Harmony evaluates the coordination and functionality of color combinations, ensuring color palette choices balance aesthetics with effective information communication (Harrower and Brewer, 2003; Gramazio et al., 2017).
@@
Adaptive Question Generation Mechanism. Based on the evaluation framework, we have devised an adaptive question generation process (detailed workflow shown in Figure 3). This process begins by leveraging GPT-4o to extract metadata from the chart, such as its type and visual elements. Subsequently, it rewrites questions by populating predefined templates based on this metadata, generating highly customized questions and their corresponding five-point scoring criteria for the six evaluation sub-dimensions. For instance, under the data fidelity dimension, the generated question specifically targets the chart’s data presentation accuracy, asking whether the “horizontal bar chart accurately represents injury counts with a proper zero baseline and proportional bar lengths.” The scoring criteria focuses on visually detectable issues: a score of 1 identifies “major distortions, such as truncated axes or misleading scales that exaggerate differences,” whereas a score of 5 confirms a “highly faithful representation where bar lengths are strictly proportional to the displayed values.” This approach ensures that the evaluation questions are closely aligned with the specific visualization content. For more detailed examples, please refer to Appendix C.1.
@@ 3.1.3 Expert Annotation and Quality Control
Stage 1: Initial Annotation. We recruited 603 highly qualified crowdsourcing workers through the CloudResearch platform (CloudResearch, 2022). To ensure annotation quality, we set strict screening criteria (see Appendix A.3.1 for details) and designed a dedicated annotation interface (Appendix A.3.3). Crucially, we embedded validation checks to identify and filter out inattentive responses (examples in Appendix A.3.4). Each of the 3,090 samples was scored by three independent annotators across six evaluation dimensions (task design details in Appendix A.3.2), generating an initial scoring matrix.
Stage 2: Quality Control. To address scoring disagreements among annotators, we designed a systematic conflict identification and resolution mechanism based on established crowdsourcing quality control and statistical evaluation theory (Gadiraju et al., 2015; Rousseeuw and Leroy, 2005; Brennan, 2001). The system first identifies high-disagreement samples by analyzing score variance, then algorithmically generates candidate resolution strategies including outlier removal, malicious scoring detection, and sub-dimensional bias correction. These algorithm-generated suggestions are processed and ranked before being submitted to the expert team for final review (complete algorithmic details and parameters in Appendix A.3.4).
Stage 3: Expert Validation. Three experts with visualization analysis experience independently reviewed all samples using a dedicated interface (Appendix A.3.5), selecting, modifying, or rejecting algorithm-generated candidate solutions. For complex cases, the team reached consensus through discussion. Through this rigorous process, we built a high-quality human scoring benchmark for all 3,090 samples as the gold standard for model evaluation. Detailed annotation quality analysis is provided in Appendix A.4.
@@ 4 VisJudge: A Specialized Model for Visualization Evaluation
The primary goal of VisJudge-Bench is to systematically evaluate MLLMs’ capabilities in visualization aesthetics and quality assessment. Building on this benchmark, we explore a key question: can domain-specific fine-tuning improve models’ visualization evaluation capabilities? To answer this question, we develop VisJudge by fine-tuning multiple open-source multimodal architectures at different parameter scales. In Section 5, we conduct unified evaluation of these fine-tuned models alongside closed-source and open-source baselines on the same test set.
@@
Model Training. We fine-tuned four representative open-source multimodal models as base models: Qwen2.5-VL-7B-Instruct and Qwen2.5-VL-3B-Instruct (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025), and Llava-v1.6-mistral-7B (Liu et al., 2024). These models span different architectures and parameter scales (3B to 8B), enabling comprehensive evaluation of our training approach’s cross-architecture generalization. All models are trained to generate quality scores (1.0-5.0) and rationales aligned with human expert judgments. We employ reinforcement learning with the GRPO algorithm (Shao et al., 2024), using a composite reward function combining accuracy reward (minimizing prediction error) and format reward (ensuring structured outputs) (Shi et al., 2025; Wu et al., 2025). Formal reward definitions are detailed in Appendix D.2. For parameter-efficient fine-tuning, we adopted Low-Rank Adaptation (LoRA) (Hu et al., 2022). All models were trained using consistent hyperparameter configurations across architectures. Training used 5 epochs with a learning rate of 1e-5. Detailed configurations are in Appendix D.3.
@@ 5.1 Experimental Settings
Evaluation Setup. We evaluate 12 representative MLLMs: GPT-5, GPT-4o, Claude-4-Sonnet, Claude-3.5-Sonnet, Gemini-2.0-Flash, Gemini-2.5-Pro, Qwen2.5-VL (3B/7B/32B/72B-Instruct) (Bai et al., 2025), InternVL3-8B (Zhu et al., 2025), Llava-v1.6-mistral-7B (Liu et al., 2024), and their corresponding fine-tuned variants, which we collectively refer to as VisJudge on a balanced test set of 648 samples (see Appendix A.2.7 for distribution details). Each model provides 1-to-5 scores with justifications based on our evaluation framework. Following human annotation procedures, we run each model three times and average the results. All inference uses vLLM on four NVIDIA A6000 (48GB) GPUs with bfloat16 precision and a temperature of 0.8.
@@ 5.2 Experimental Results and Analysis
Metric Type Model Overall Fidelity Expressiveness Aesthetics
Readability Insight Design Style Composition Color
MAE (↓) Closed- source Claude-3.5-Sonnet 0.824 0.978 0.904 1.154 0.783 0.940 0.862
Claude-4-Sonnet 0.622 0.841 0.757 0.832 0.679 0.734 0.785
Gemini-2.0-Flash 0.682 0.829 0.912 0.820 0.638 0.729 0.798
Gemini-2.5-Pro 0.662 1.243 0.945 0.900 0.840 0.918 0.980
GPT-4o 0.610 0.988 0.806 0.744 0.609 0.695 0.657
GPT-5 0.553 0.862 0.781 0.778 0.649 0.699 0.682
Open- source Qwen2.5-VL-3B-Instruct 0.821 1.085 1.258 1.088 0.723 0.727 0.808
Qwen2.5-VL-7B-Instruct 0.847 1.171 1.296 0.858 0.756 0.812 0.772
Qwen2.5-VL-32B-Instruct 0.703 0.909 0.987 0.801 0.678 0.761 0.719
Qwen2.5-VL-72B-Instruct 0.702 0.999 0.926 0.811 0.663 0.735 0.717
InternVL3-8B 0.793 1.234 1.193 0.870 0.679 0.759 0.753
Llava-v1.6-mistral-7B 0.724 0.929 1.124 0.939 0.801 0.814 0.818
VisJudge (Ours) InternVL3-8B 0.541 0.797 0.769 0.691 0.608 0.645 0.616
Llava-v1.6-mistral-7B 0.496 0.767 0.695 0.775 0.643 0.574 0.598
Qwen2.5-VL-3B-Instruct 0.491 0.705 0.721 0.696 0.625 0.571 0.616
Qwen2.5-VL-7B-Instruct 0.421 0.661 0.648 0.677 0.580 0.545 0.604
MSE (↓) Closed- source Claude-3.5-Sonnet 1.009 1.577 1.306 1.989 0.994 1.467 1.198
Claude-4-Sonnet 0.603 1.182 0.973 1.147 0.774 0.934 1.037
Gemini-2.0-Flash 0.720 1.182 1.326 1.120 0.673 0.925 1.057
Gemini-2.5-Pro 0.677 2.291 1.478 1.371 1.112 1.325 1.460
GPT-4o 0.577 1.562 1.063 0.921 0.627 0.823 0.729
GPT-5 0.486 1.219 0.989 0.970 0.720 0.862 0.810
Open- source Qwen2.5-VL-3B-Instruct 1.028 1.872 2.400 1.852 0.868 0.903 1.046
Qwen2.5-VL-7B-Instruct 1.045 2.051 2.413 1.177 0.938 1.093 0.996
Qwen2.5-VL-32B-Instruct 0.756 1.361 1.528 1.067 0.728 0.988 0.880
Qwen2.5-VL-72B-Instruct 0.762 1.626 1.373 1.092 0.713 0.932 0.870
InternVL3-8B 0.934 2.240 2.088 1.219 0.747 0.981 0.968
Llava-v1.6-mistral-7B 0.817 1.513 1.960 1.403 1.007 1.071 1.076
VisJudge (Ours) InternVL3-8B 0.437 1.068 0.951 0.822 0.638 0.736 0.655
Llava-v1.6-mistral-7B 0.383 0.919 0.752 0.952 0.642 0.541 0.619
Qwen2.5-VL-3B-Instruct 0.377 0.855 0.850 0.822 0.620 0.539 0.603
Qwen2.5-VL-7B-Instruct 0.286 0.747 0.690 0.756 0.545 0.498 0.578
Corr. (↑) Closed- source Claude-3.5-Sonnet 0.395 0.325 0.492 0.365 0.455 0.137 0.259
Claude-4-Sonnet 0.465 0.393 0.550 0.452 0.421 0.163 0.228
Gemini-2.0-Flash 0.395 0.372 0.459 0.417 0.459 0.157 0.209
Gemini-2.5-Pro 0.265 0.178 0.379 0.353 0.445 0.193 0.208
GPT-4o 0.482 0.381 0.539 0.442 0.471 0.277 0.363
GPT-5 0.428 0.255 0.439 0.382 0.463 0.276 0.295
Open- source Qwen2.5-VL-3B-Instruct 0.272 0.199 0.222 0.275 0.338 0.130 0.155
Qwen2.5-VL-7B-Instruct 0.341 0.341 0.352 0.281 0.357 0.149 0.155
Qwen2.5-VL-32B-Instruct 0.435 0.348 0.468 0.408 0.449 0.200 0.268
Qwen2.5-VL-72B-Instruct 0.440 0.331 0.479 0.416 0.435 0.165 0.251
InternVL3-8B 0.409 0.323 0.407 0.344 0.419 0.216 0.170
Llava-v1.6-mistral-7B 0.180 0.201 0.137 0.160 0.188 0.064 0.076
VisJudge (Ours) InternVL3-8B 0.660 0.533 0.594 0.545 0.499 0.391 0.420
Llava-v1.6-mistral-7B 0.605 0.226 0.536 0.443 0.432 0.406 0.403
Qwen2.5-VL-3B-Instruct 0.648 0.533 0.581 0.579 0.504 0.490 0.402
Qwen2.5-VL-7B-Instruct 0.687 0.574 0.628 0.576 0.568 0.513 0.385